# Enhancing Robot Navigation Policies with Task-Specific Uncertainty Management

### v2

@@ Appendix A Additional Experimental Results
Appendix B Implementation Details
B.1. Training the TSUM Generation Model
In this section, we provide a detailed description of the dataset used to train the embedding functions and attention networks for the Task-Specific Uncertainty Map (TSUM) generation model. We explain the data collection process, the nature of the tasks included, and how the data was annotated and processed to facilitate effective learning.
B.1.1. Dataset Description and Collection Process
The dataset consists of paired semantic concepts and spatial locations with known relationships, collected using an ASV in a controlled aquatic environment. We used this data alongside the open-sourced Virtual Ocean Robotics Challenge simulator [^3] https://github.com/osrf/vorc. The ASV was equipped with a GPS, an IMU, and a YSI EXO2 sonde for measuring water quality parameters. The environment featured a variety of navigational elements such as docks, fountains, and virtual designated zones to provide contextual information.
To construct the dataset, we designed and executed a series of navigation tasks that encompassed diverse operational scenarios. During each task, both in simulation and on hardware, sensor data, positional information, and timestamped actions were recorded. Approximately 50 natural language task descriptions were hand-crafted with placeholders to include specific semantic elements that could be directly associated with spatial locations within the environment. The dataset includes over 8,000 task instances, ensuring comprehensive coverage of different task types and constraints.
B.1.2. Task Categories and Examples
@@
B.1.3. Natural Language Processing and Embedding Generation
@@
B.1.4. Data Annotation and Processing
For each task, the natural language descriptions were parsed to extract semantic concepts such as goals, landmarks, actions, and constraints. These concepts were mapped to spatial coordinates using the ASV’s GPS data and predefined environmental maps. The pairing of semantic elements with precise spatial locations enabled the creation of meaningful training examples for the embedding functions.
Positive pairs (related semantic concepts and spatial locations) and negative pairs (unrelated concepts and locations) were included to optimize the contrastive loss function during training. This ensures that the model learns to distinguish between relevant and irrelevant semantic-spatial relationships, enhancing its ability to generate accurate and contextually appropriate uncertainty maps.
B.2. Hyperparameters
@@
Hyperparameter Name Brief Explanation Value $w_{\Phi}$ Weight for task relevance function $\Phi^{\tau}(l)$ 0.5 $w_{\mathcal{C}}$ Weight for task-specific constraints $\mathcal{C}^{\tau}(l)$ 0.3 $w_{\mathcal{E}}$ Weight for environmental factors $\mathcal{E}(l)$ 0.2 Embedding dimension $n$ Dimensionality of semantic and spatial embedding spaces 256 Margin parameter $\delta$ Margin in triplet loss function 0.5 Balancing parameter $\lambda$ Balances triplet and attention losses in total loss function 1.0 Discount factor $\gamma$ Discount factor in reinforcement learning objective 0.99 Initial temperature $\alpha$ Initial temperature parameter in SAC 0.2 Target entropy $\bar{\mathcal{H}}$ Target entropy for policy entropy term $-2$ Soft target update coefficient $\tau$ Update rate for target Q-networks $Q_{\bar{\phi}_{i}}$ 0.005 Replay buffer size $|\mathcal{D}|$ Capacity of experience replay buffer $1\times 10^{6}$ Batch size Number of samples per training batch 256 Learning rate for policy network $\pi_{\theta}$ Learning rate for policy network parameters $3\times 10^{-4}$ Learning rate for Q-value networks $Q_{\phi_{i}}$ Learning rate for Q-value network parameters $3\times 10^{-4}$ Learning rate for temperature $\alpha$ Learning rate for adjusting temperature parameter $3\times 10^{-4}$ Policy network architecture Neural network structure for $\pi_{\theta}$  Number of hidden layers Number of layers in policy network 2  Hidden units per layer Units in each hidden layer of policy network 256  Activation function Activation function used in policy network ReLU Q-value networks architecture Neural network structure for $Q_{\phi_{i}}$  Number of hidden layers Number of layers in Q-value networks 2  Hidden units per layer Units in each hidden layer of Q-value networks 256  Activation function Activation function used in Q-value networks ReLU Maximum steps per episode Max steps allowed per episode 1000
Table 2. Hyperparameters for GUIDEd-SAC.
@@
Table 3. Hyperparameters for the baseline and abalation methods. Unless specified, methods use the same hyperparameters as in GUIDEd SAC (see Table 2). Only differences from GUIDEd SAC are listed for each method.

### v3

@@ VII Additional Experimental Results
VII-A Implementation Details
VII-A1 Task Categories and Examples
@@
VII-A2 Natural Language Processing and Embedding Generation
@@
VII-A3 Dataset and Data Processing
The dataset used for training our TSUM generation model consists of overhead imagery collected from various marine environments and simulation scenes, carefully curated to capture a wide range of operational scenarios. The data collection process spanned multiple weather conditions to ensure robustness and generalizability of the trained model.
For the labeling process, we employed experts with experience in autonomous navigation. These experts followed an annotation protocol to maintain consistency across the dataset. The protocol involved examining each 224×224 pixel patch of overhead imagery and assigning binary labels indicating relevance to predefined subtasks and constraints. The labeling criteria were established through iterative refinement to ensure practical applicability.
To address class imbalance issues in our dataset, we implemented a comprehensive sampling strategy. Many critical navigation scenarios, such as complex docking maneuvers or navigation around certain obstacles, were naturally underrepresented in the raw data collection. We addressed this through strategic oversampling of these crucial but rare scenarios. Specifically, patches containing these underrepresented classes were duplicated in the training set, while maintaining cross-validation splits to prevent overfitting. For negative examples, we employed a systematic sampling approach, randomly selecting patches that domain experts had deemed irrelevant to specific subtasks or constraints, while ensuring a balanced representation across different environmental conditions.
The environmental variations in our dataset were curated to capture the diverse conditions encountered in real-world marine operations. Our collection includes patches from both real-world lake environments and high-fidelity simulated scenarios. The real-world data encompasses variations in weather conditions (clear, overcast, and light rain), lighting conditions (morning, midday, and evening), and seasonal changes affecting water conditions. Shoreline features vary from natural boundaries to man-made structures, and obstacle density ranges from sparse open waters to crowded marina environments. The simulated data complements these real-world scenarios by providing additional coverage of rare but critical scenarios that are difficult to capture in real-world data collection.
For data storage and organization, we implemented a metadata management system. Each image patch is stored in a georeferenced index using a custom CSV format, which maintains associations between image patches, their geographical coordinates, and corresponding labels. This schema ensures efficient retrieval during training and enables seamless integration with existing robotic navigation systems. The metadata includes essential information such as timestamp, environmental conditions, and annotation confidence scores from domain experts.
Our data augmentation pipeline was designed to enhance model robustness while preserving critical visual cues for marine navigation. The primary augmentations include random rotations (in 90-degree increments) and horizontal flips, which reflect the rotational invariance of navigation tasks while maintaining the natural appearance of water features. Notably, we deliberately avoided color jittering and intensity transformations, as these could distort important water-related visual cues that are crucial for accurate navigation. The augmentation parameters were tuned through extensive experimentation to find the optimal balance between increasing data diversity and maintaining task-relevant features. All image patches undergo consistent preprocessing to ensure uniformity in the training data. This includes standardization to zero mean and unit variance, computed across the training set, and resizing to maintain consistent spatial resolution across different source imagery. The preprocessing pipeline also includes automated quality checks to identify and filter out patches with excessive noise, sensor artifacts, or poor visibility conditions that could negatively impact model training.
VII-A4 Technical Infrastructure and Hardware Setup
Our technical infrastructure was designed to efficiently handle the computational demands of processing large-scale overhead imagery and training complex neural networks. The foundation of our data processing pipeline centers on the conversion of geographical coordinates to discrete 224×224 pixel patches, which serve as the basic unit of analysis for our TSUM generation system. The coordinate-to-patch mapping system employs a gridding algorithm that subdivides large overhead maps into a regular grid of fixed-size patches. Each patch is assigned center coordinates that correspond to real-world geographical positions. The mapping process utilizes a custom spatial indexing structure that enables efficient lookup of relevant patches given any arbitrary location in the operational space. To handle edge cases where locations fall between patch boundaries, we implemented a nearest-center assignment strategy that ensures every point in the operational space maps to exactly one patch while maintaining spatial continuity.
To maximize processing efficiency, we developed a parallelized data generation pipeline. The system utilizes a custom-built multithreaded tiling script that leverages all available CPU cores to concurrently process large imagery datasets. This parallelization is implemented using Python’s multiprocessing library, with careful attention to memory management to prevent resource exhaustion when handling particularly large maps. The tiling process is coordinated by a master thread that manages work distribution and ensures balanced load across all available processors. Each worker thread independently processes assigned regions of the input imagery, generating patches and associated metadata in parallel.
The runtime environment is built on Ubuntu 20.04 LTS with CUDA 11.8 and cuDNN 8.7, optimized for deep learning workloads. We utilize Docker containers to ensure consistency across different compute nodes and to simplify deployment. The container images are based on NVIDIA’s NGC PyTorch container, customized with additional dependencies required for our specific workload. This containerized approach ensures reproducibility and enables easy scaling across different hardware configurations. Data movement between storage and compute nodes is optimized using a custom data loading pipeline built on top of PyTorch’s DataLoader class. We implemented prefetching mechanisms that load and preprocess data for upcoming batches while the current batch is being processed on the GPUs. This approach effectively hides I/O latency and ensures near-continuous GPU utilization. The data loading pipeline includes automatic checkpointing to enable recovery from system failures without losing progress.
The typical processing workflow for a complete dataset involves approximately 2-3 hours of initial data tiling and preprocessing, followed by 10 epochs of model training. The training process is distributed using PyTorch’s DistributedDataParallel, with gradient synchronization optimized for our specific network architecture and batch sizes. Our implementation achieves approximately 71% GPU utilization during training, with the remaining overhead primarily attributed to necessary data loading operations.
Resource monitoring and system health checks are performed using a combination of Prometheus for metrics collection and Grafana for visualization. This monitoring infrastructure allows us to track system performance, identify bottlenecks, and optimize resource utilization in real-time. Additionally, we maintain comprehensive logs of all processing steps, enabling detailed analysis of system performance and facilitating debugging when necessary.
A crucial component of our technical infrastructure is our high-fidelity simulation environment, built using Unity3D (2022.3.16f1) in conjunction with ROS2 Humble. This simulator serves as a vital tool for both data generation and policy validation. The Unity environment provides physically accurate water dynamics for fluid simulation, capable of modeling complex wave patterns, water resistance, and hydrodynamic forces. We implemented custom shaders to accurately render water surface properties, ensuring visual fidelity crucial for training vision-based navigation systems.
The simulator is tightly integrated with ROS2 through a custom bridge that enables bidirectional communication between Unity and ROS2 nodes. This integration allows for seamless transfer of sensor data, control commands, and state information. We implemented the full suite of sensors found on our physical ASV, including simulated GPS (with configurable noise patterns), IMU (with drift characteristics matching real hardware), and cameras (with accurate lens distortion and environmental effects). The simulator also includes detailed models of environmental factors such as wind effects, current patterns, and varying lighting conditions.
For training data generation, the simulator can be run in a distributed headless manner across instances, each generating different scenarios and environmental conditions. We used an off-shelf scenario generation system that creates diverse training situations, including varying obstacle configurations, weather conditions, and traffic patterns. The simulator supports both synchronous and asynchronous operation modes, allowing for rapid data collection when generating training datasets and real-time operation when validating policies.
VII-B Hyperparameters
@@
Hyperparameter Name Brief Explanation Value TSUM Aggregation $w_{\Phi}$ Weight for subtask relevance function $\Phi^{\tau}(l)$ 0.5 $w_{\mathcal{C}}$ Weight for constraint relevance $\mathcal{C}^{\tau}(l)$ 0.3 $w_{\mathcal{E}}$ Weight for environmental factors $\mathcal{E}(l)$ 0.2 CLIP Fine-Tuning CLIP embedding dimension $d$ Output dimension of ViT-B/32 encoders (text/image) 512 Number of frozen layers Initial layers in CLIP left frozen to retain general cross-modal knowledge 6 Fine-tuning epochs Epochs used to adapt the unfrozen CLIP layers 10 Learning rate for CLIP LR for training the unfrozen CLIP layers $1\times 10^{-5}$ Alignment weighting $\lambda_{\text{align}}$ Weight on additional alignment loss $\mathcal{L}_{\text{align}}$ 1.0 Contrastive temperature $\zeta$ Temperature hyperparameter for CLIP contrastive loss 0.07 Reinforcement Learning (SAC) Discount factor $\gamma$ Discount factor in the RL objective 0.99 Initial temperature $\alpha$ Initial entropy temperature in SAC 0.2 Target entropy $\bar{\mathcal{H}}$ Target policy entropy term in SAC $-2$ Soft target update $\tau$ Polyak averaging coefficient for target Q-networks 0.005 Replay buffer size $|\mathcal{D}|$ Maximum capacity of the experience replay buffer $1\times 10^{6}$ Batch size Number of samples per RL training batch 256 Learning rate for policy network LR for optimizing the policy $\pi_{\theta}$ $3\times 10^{-4}$ Learning rate for Q-value networks LR for optimizing the Q-value functions $Q_{\phi_{i}}$ $3\times 10^{-4}$ Learning rate for temperature $\alpha$ LR for adjusting the SAC temperature $3\times 10^{-4}$ Policy network architecture MLP for $\pi_{\theta}$ with hidden layers and activation 2 layers, 256 units each, ReLU Q-value networks architecture MLP for $Q_{\phi_{i}}$, identical structure 2 layers, 256 units each, ReLU Maximum steps per episode Upper bound on steps in each training episode 1000
TABLE II: Hyperparameters for the CLIP-based TSUM generation and SAC policy learning.
@@
TABLE III: Hyperparameters for the baseline and abalation methods. Unless specified, methods use the same hyperparameters as in GUIDEd SAC (see Table II). Only differences from GUIDEd SAC are listed for each method.